Of all possible quotients of the numbers 10, -1/2, -5, and 1/5 which tow produce the largest quotient? Which two produce the smallest?
Answer by checkley71(8403)

-1/2*-5=5/2=2.5 the largest quotient.
10*-5=-50 is the smallest quotient.